OpenStudy (anonymous):

Find cos2 theta given the cos is 4/5

OpenStudy (anonymous):

I have no idea how to do this question. Would you just multiply it That would give you 16/ 25

OpenStudy (anonymous):

@diamolnd2 do you mean \[\cos^{2} \theta?\] \[\cos^{2} \theta = (\cos \theta)^{2}\]

OpenStudy (anonymous):

no its not squared

OpenStudy (anonymous):

sin2\[\theta \] + cos\[\theta \] = 0

OpenStudy (anonymous):

sorry thats a different problem, im confused on that one two though

OpenStudy (anonymous):

if it's not squared, do you mean \[\cos 2\theta?\] you can use this formula: \[\cos 2\theta=2\cos^{2} \theta-1\]

OpenStudy (anonymous):

yes that was what i ment

OpenStudy (anonymous):

2(16/25)-1 =7/25